# BioMoDiffuse: Physics-Guided Biomechanical Diffusion for Controllable and Authentic Human Motion Synthesis

### v1

@@ Appendix B Implementation Details
For training, we employ DDPM [20] with 1,000 denoising steps ($T=1,000$) and linearly varying variances $\beta_{t}$ from 0.0001 to 0.02 in the forward process. The model is trained using AdamW [40] with an initial learning rate of 0.0002 and a weight decay of 0.01 for 50,000 iterations during pre-training, followed by 10,000 iterations of fine-tuning with a batch size of 64 on an NVIDIA A40 GPU. The learning rate is reduced by a factor of 0.9 every 5,000 steps. Gradient norms are clipped to a maximum value of 1 during gradient descent. For class-conditional generation (CFG) [19], text conditioning $c$ is set to $\emptyset$ for 10% of the samples, and the physical network is ignored with a probability of $P_{c}=10\%$. Additionally, we apply Exponential Moving Average (EMA) to smooth the parameter updates and reduce mutations and oscillations.
For inference, we utilize the SDE variant of the second-order DPM-Solver++ [41] with Karras sigmas [28] for 10-step sampling. The scale for text conditioning CFG is set to 2.5.
@@
In our implementation of pose refinement, we used WHAM [60] as the baseline. However, the official WHAM code repository only provides versions trained on five datasets: AMASS [44], InstaVariety [27], MPI-INF-3DHP [45], Human3.6M [23], and 3DPW [70]. It does not include a version that incorporates the BEDLAM [4] dataset for training. As a result, the baseline results in our work differ from those presented in the WHAM paper.
Appendix C Network Architecture
Table 7: Model Architecture of BioMoDiffuse on dataset HumanML3D.
Model Module Details
BioMo Diffuse Physical Network (1-1)
• GRU layer
• AccelNet:
   - Motion MLP
   - Muscle MLP
   - Physics Decoder
   - Muscle Decoder
   - M_Net
   - F_Net
   - C_Net
   - Pose2EMG
Cond Unet1D (1-2)
• Time MLP
• Downsampling Blocks (×4) :
   - Conv1D
   - Cross Attention
   - Residual Blocks
• Mid Blocks (×2) :
• Upsampling Blocks (×4) :
   - Transposed Conv1D
   - Residual Blocks
Tab. 7 illustrates the network architecture of BioMoDiffuse trained on HumanML3D [14]. Compared to the baseline UNet, the addition of our physical network only increases the parameter count by $6.68\%$. However, this results in a significant improvement in the FID score, which decreases by $48.8\%$ (from 0.129 to 0.068). Other metrics also show notable improvements, demonstrating that the proposed physical network is both lightweight and effective.
Appendix D Derivation of Euler-Lagrange equation
The classical Euler-Lagrange equation is formulated as:
$\frac{d}{dt}\left(\frac{\partial L}{\partial\dot{q}}\right)-\frac{\partial L}{\partial q}=Q,$ (20)
where $L=T-V$ is the Lagrangian, $T$ denotes kinetic energy, $V$ represents potential energy, $q$ are generalized coordinates, and $Q$ encompasses non-conservative forces such as joint torques and external contact forces.
By modeling the human body as a rigid-body system, its kinetic energy can be expressed as:
$T=\frac{1}{2}\dot{q}^{T}M(q)\dot{q},$ (21)
where $M(q)$ is the inertia matrix. The potential energy $V(q)$ is defined by the gravitational field. Substituting $L=T-V$ into the Euler-Lagrange equation yields:
$\frac{d}{dt}\left(\frac{\partial T}{\partial\dot{q}}\right)-\frac{\partial T}{\partial q}+\frac{\partial V}{\partial q}=Q.$ (22)
Substituting the explicit form of $T$, we derive:
$M(q)\ddot{q}+\dot{M}(q)\dot{q}-\frac{1}{2}\dot{q}^{T}\frac{\partial M}{\partial q}\dot{q}+G(q)=Q,$ (23)
where $G(q)=\frac{\partial V}{\partial q}$. Leveraging the symmetry of $M(q)$, we reorganize indices and employ Christoffel symbols to define the Coriolis matrix:
$C(q,\dot{q})\dot{q}=\sum_{i,j}\frac{1}{2}\left(\frac{\partial M_{kj}}{\partial q_{i}}+\frac{\partial M_{ki}}{\partial q_{j}}-\frac{\partial M_{ij}}{\partial q_{k}}\right)\dot{q}_{i}\dot{q}_{j}.$ (24)
It can be rigorously shown that $C(q,\dot{q})\dot{q}=\dot{M}(q)\dot{q}-\frac{1}{2}\dot{q}^{T}\frac{\partial M}{\partial q}$, demonstrating that the Coriolis term arises from the derivative of the inertia matrix and quadratic velocity coupling, which depend solely on the body’s pose.
Figure 6: More samples from our BioMoDiffuse for text-to-motion synthesis. Samples generated with text prompts of HumanML3D’s test set. We recommend the supplemental video to see more motion results.
Appendix E Comparison on Samplers
Table 8: Comparative Evaluation of Motion Generation Methods. We compare different methods on generation quality, efficiency, and physical plausibility metrics using the HumanML3D dataset.
[…]
GT - - $0.002$ $2.974$ $0.797$ $0.124$ $0.682$ $0.358$ $2.60$ $0.205$ $0.057$ $0.000$ $0.000$
DEIS $10$ $3.659$ $0.148$ $2.732$ $0.843$ $0.298$ $2.388$ $0.380$ $0.91$ $0.195$ $0.054$ $0.013$ $0.002$
DEIS w/o Karras $10$ $3.615$ $0.146$ $2.723$ $0.844$ $0.271$ $2.285$ $0.381$ $0.88$ $0.195$ $0.051$ $0.013$ $0.002$
DPMSolver++ w/o Karras $10$ $3.664$ $0.126$ $2.755$ $0.840$ $0.147$ $1.071$ $0.374$ $1.39$ $0.196$ $0.050$ $0.012$ $0.001$
DPMSolver++ (20) $20$ $7.610$ $0.090$ $2.781$ $0.834$ $0.143$ $0.969$ $0.382$ $1.44$ $0.203$ $0.059$ $0.013$ $0.003$
DPMSolver++ (10) $10$ $2.553$ $0.071$ $2.784$ $0.835$ $0.135$ $0.959$ $0.385$ $1.45$ $0.204$ $0.056$ $0.013$ $0.002$
DDIM $50$ $19.82$ $0.123$ $2.741$ $0.839$ $0.132$ $0.933$ $0.368$ $1.57$ $0.185$ $0.054$ $0.012$ $0.002$
DDPM $1000$ $290.30$ $0.088$ $2.771$ $0.833$ $0.147$ $0.981$ $0.385$ $1.43$ $0.209$ $0.060$ $0.013$ $0.003$
In this section, we present the impact of different samplers on the generation performance of BioMoDiffuse. We evaluated the following samplers on our model: DDPM [20] , DDIM [62], DPMSolver [41], and DEIS [84]. Additionally, we incorporated Karras Sigma into DEIS and DPMSolver. Karras Sigma corresponds to constant-velocity thermal diffusion, which has been shown to ensure faster and higher-quality generation in the field of image synthesis. As shown in Tab. 8, we observe that DEIS fails to generate sufficiently smooth motions, while among the other samplers, DPMSolver achieves superior results in terms of FID and physical plausibility metrics. Furthermore, the inclusion of Karras Sigma contributes to improved FID scores. Based on these findings, we ultimately selected DPMSolver with Karras Sigma as the preferred configuration for our framework.
Appendix F User Study
Figure 7: User study interface for evaluating StableMoFusion and our model on HumanML3D.
In the user study, we generated multiple samples using the evaluation code on the HumanML3D test set, filtering out semantically redundant examples to create a final set of 20 distinct samples. Participants were asked to rate the generated motions on a 1–5 scale across four dimensions: ”Generation Quality,” ”Motion Smoothness,” ”Physical Plausibility,” and ”Text Alignment.” We invited 30 users to compare the motion generation results of our model, StableMoFusion, and Ground Truth based on these 20 prompts. To ensure unbiased feedback, the samples were presented in random order, preventing participants from identifying which motions were generated by our model.
Appendix G More Qualitative Results
Fig. 6 demonstrates the generation performance of our model across a broader set of prompts, while Fig. 8 provides a comparative analysis of our model against state-of-the-art methods on an extended set of prompts. The results indicate that our model consistently achieves superior performance across a wider range of motion types.
Figure 8: The comparison of the state-of-the-art methods on text-guided motion synthesis task. Samples generated with text prompts of HumanML3D’s test set. We recommend the supplemental video to see more motion results.

### v2

@@ IV-F Implementation Details
All experiments were conducted on a single NVIDIA A40 GPU. For our implementation of motion generation, we adopt StableMoFusion as the baseline. We employ DDPM [39] with 1,000 denoising steps ($T=1{,}000$) and linearly varying variances $\beta_{t}$ from 0.0001 to 0.02 in the forward process. The model is trained using AdamW [71] with an initial learning rate of 0.0002 and a weight decay of 0.01 for 50,000 iterations during pre-training, followed by 10,000 iterations of fine-tuning with a batch size of 64. Additionally, we apply an Exponential Moving Average (EMA) to smooth parameter updates and reduce instability and oscillations. During inference, we utilize the SDE variant of the second-order DPM-Solver++ [72] with Karras sigmas [73] for 10-step sampling. We set the CFG scale for text conditioning to 2.5.
@@
For our implementation of pose refinement, we adopted WHAM [4] as the baseline. However, the official WHAM code repository provides models trained only on five datasets: AMASS [70], InstaVariety [74], MPI-INF-3DHP [75], Human3.6M [76], and 3DPW [17], and does not include a version trained with the BEDLAM [77] dataset. Consequently, the baseline results reported in our work differ slightly from those presented in the original WHAM paper.
Figure 8: Editing generated human motions to different movement amplitudes. By manually modifying the acceleration values in BioVAE, controllable motion generation can be achieved. Adjusting the acceleration for a single joint or all joints allows the generated motions to exhibit varying amplitudes.
V Conclusion
In this paper, we propose BioVAE, a biomechanics-aware framework designed for a wide range of tasks, including motion generation, pose estimation, motion refinement, and motion editing. BioVAE is inspired by extended Euler-Lagrange equations and achieves deep integration with the diffusion process through a carefully designed loss function. Furthermore, it incorporates joint learning from electromyographic (EMG) signals and acceleration data. Experimental results demonstrate that our approach outperforms previous models in both generation quality and physical plausibility. However, data limitations have so far prevented us from evaluating our approach on high-frame-rate datasets with richer annotations. Addressing this limitation in future work could enable BioVAE to deliver even stronger results for practical industrial applications.